OpenStudy (anonymous):

Simplify 15-3d/2d-10 HELP!!

OpenStudy (anonymous):

3(5-d)/2(d-5)

OpenStudy (anonymous):

i need help determining the answer can it be simplified further?

OpenStudy (anonymous):

@laxman03

OpenStudy (anonymous):

can someone please help me with the answer.....

OpenStudy (anonymous):

is there a solution?

zepdrix (zepdrix):

\[\large \frac{15-3d}{2d-10}\]Factoring 3 out of each term in the top, 2 out of each term in the bottom, gives us,\[\large \frac{3(5-d)}{2(d-5)}\]From here, let's factor a -1 out of each term in the top,\[\large \frac{-3(d-5)}{2(d-5)}\]Was that step confusing at all?

OpenStudy (anonymous):

jeez wow i didnt know you could do that.......

zepdrix (zepdrix):

lol ya math is cool XD

OpenStudy (anonymous):

ur the best finally a solution!

zepdrix (zepdrix):

So you see the cancellation that comes next? :)

OpenStudy (anonymous):

yep -3/2

zepdrix (zepdrix):

ok cool